OpenStudy (anonymous):

Of those invited to a wedding, 2/7 were friends of the bride. The friends of the groom numbered 65. How many people were invited to the wedding

OpenStudy (kropot72):

The 65 friends of the groom are (1 - 2/7) = 5/7 of the people at the wedding. Therefore the number of people at the wedding is found from: \[\frac{65\times 7}{5}=?\]

OpenStudy (anonymous):

91?

OpenStudy (kropot72):

Good! There were 91 people at the wedding.

OpenStudy (anonymous):

awesome tyvm!

OpenStudy (kropot72):

You're welcome :)